OpenStudy (anonymous):

Which term is -401 in the arithmetic sequence -5, -9, -13…? A. 56th B. 100th C. 101st D. 109th

OpenStudy (gudden):

Well find the common difference of the sequence i.e. d by subtracting two terms -9 -(-5) = - 4 and first term i.e. a of the sequence is -5 so, plug these in the formula Term = n/2 ( 2a + (n-1)d) -401 *2 = n(-10 + (n - 1)-4 => -802 = n( -10 -4n +4 ) => -802 = n( -6 - 4n) => -802 = -6n - 4n^2 => 4n^2 + 6n - 802 = 0 solve this quadratic equation.... and you would get two roots... Eliminate the negative root and +ve root would be your answer... Good luck